THE PRODUCT OF TWO CONSECUTIVE POSITIVE EVEN INTEGERS IS 440?
Answer by checkley75(3666)

x*(x+2)=440
x^2+2x-440=0
(x-20)(x+22)=0
x-20=0
x=20 answer.
x+2=22 answer.
proof
20*22=440
440=440